OpenStudy (anonymous):

I really need help! i have to make three equations:One function, f(x), with two real rational solutions. One function, g(x), with two real irrational solutions. One function, h(x), with two complex solutions.

OpenStudy (ranga):

Do you know what the discriminant of a quadratic equation is?

OpenStudy (anonymous):

no

OpenStudy (ranga):

If \[\large ax ^{2} + bx + c = 0\]is the quadratic equation then the discriminant D = \[\Large D = b^{2} - 4ac\] If D is positive you will have two real roots If D is zero you will have two equal roots If D is negative you will have two complex roots If D is positive and is a perfect square you will get two real rational roots If D is positive and is not a perfect square you will get two real irrational roots

OpenStudy (anonymous):

wait is it where the number is the same?

OpenStudy (anonymous):

ok thank you

OpenStudy (ranga):

They want f(x) to have two real rational roots: That means D must be positive and a perfect square. The easiest way is to pick two rational roots, say 4 and 6. Then (x - 4)(x - 6) = 0 because the roots can be factored out. (x - 4)(x - 6) = x^2 - 4x - 6x + 24 = 0 x^2 - 10x + 24 So f(x) = x^2 - 10x + 24 is an example of a function with two rational toots: 4 and 6. That answers the first part.

OpenStudy (ranga):

If you calculate D for above you will see it is a perfect square. a = 1, b = -10 and c = 24 D = b^2 - 4ac = (-10)^2 - (4)(1)(24) = 100 - 96 = 4 which is a perfect square (square of 2) and so the roots (or the solutions) will be real and rational.

OpenStudy (ranga):

For g(x) they want two real irrational roots. And for that all we need to do is make the discriminant D not a perfect square. SO we can use the same a and b values from the previous question and just change c slightly. a = 1, b = -10, c = 23 (previously we had 24) D = b^2 - 4ac = (-10)^2 - 4(1)(23) = 100 - 92 = 8 and 8 is not a perfect square. So it will have two real irrational roots. So g(x) = x^2 - 10x + 23

OpenStudy (ranga):

To have two complex roots: just change c slightly again to make the discriminant negative. a = 1, b = -10, c = 26 D = (-10)^2 - (4)(1)(26) = 100 - 104 = -4 Discriminant is negative so two complex roots h(x) = x^2 - 10x + 26 So the answers are: f(x) = x^2 - 10x + 24 g(x) = x^2 - 10x + 23 h(x) = x^2 - 10x + 26

OpenStudy (anonymous):

Thank you so much!

OpenStudy (ranga):

you are welcome.